OpenStudy (anonymous):

The pmf of a random variable X is given below. x 1 2 3 4 5 p(x) .1+Ѳ .2+Ѳ .2 + 3 Ѳ .1- Ѳ .4 - 4 Ѳ In a test H_0: Ѳ = 0 versus H_1 : Ѳ = .1, H_0: is rejected if X=1 , X=2 or X=3. a) what is the probability of a type I error. b) is this a powerful test?

OpenStudy (anonymous):

go to your previous post

OpenStudy (anonymous):

P(type I error) = Probability of rejecting H0 when it's true. this would be P(X=1|H0)+P(X=2|H0)+P(X=3|H0)

OpenStudy (anonymous):

power is probability of rejecting when H0 is false. this would be P(X=1|H1)+P(X=2|H1)+P(X=3|H1)

OpenStudy (anonymous):

if you're going to repost, close out the original

OpenStudy (anonymous):

you there?